# Adam Wunderlich and Aric Sanders

### v1

@@ IV-C Correlated Samples
Lastly, we examined the impacts of I/Q measurement correlations on the expected PAPR. Specifically, Monte Carlo simulations were performed to assess PAPR of correlated samples generated by low-pass filtering I/Q WGN. This example gives insight into the case of band-limited Gaussian noise when the (one-sided) bandwidth is smaller than the Nyquist frequency, equal to one half the sampling rate.
@@
Mean PAPR estimated from $10^{4}$ independent Monte Carlo trials for each sample size is plotted in Fig. 5 (Top) together with the theoretical mean PAPR (11) for ideal I/Q WGN. In all cases, 95% confidence intervals were smaller than ±$\pm 0.025$ dB. We see that the mean PAPR generally decreased as the filter cutoff, $f_{c}$, decreased. Moreover, the mean PAPR was very close to the theoretical WGN value for $f_{c}\geq 0.25$, indicating that (11) is an accurate approximation when $0.25\leq f_{c}\leq 0.5$. Fig. 5 (Bottom) compares PDFs estimated with the KDE method described in Sec. IV-A for the $n=10^{4}$ sample size to the I/Q WGN theoretical PDF (6).
Figure 5: Correlated Noise Example. Top: Low-pass digital filter frequency responses with cutoff frequency, $f_{c}$, specified in cycles per sample. Middle: Mean PAPR. Bottom: PDFs for n=10 000.
@@ V Example Application to Spectral Analysis
Last, we demonstrate how our main result, (11), can be applied to spectral analysis of sampled I/Q signals to determine if a given frequency band contains only WGN. Specifically, we illustrate the application to spectrograms, which are widely used to visualize the frequency content of a nonstationary signal as it varies in time. A spectrogram is the magnitude squared of a sampled short-time Fourier transform (STFT), computed by applying the discrete Fourier transform (DFT) successively to shifted, windowed signal segments [11, Sec. 15.3].
@@
To illustrate how the above theoretical observation can be applied, we used a Deepwave Digital AIR-T 7201B [33] software defined radio (SDR) to record I/Q data at a sample rate of 30 Ms/s, for 30 ms at a center frequency of 3700 MHz. A general-purpose antenna, band-pass filter, and low-noise amplifier were used as an RF front-end prior to the SDR to amplify signals in the 3 GHz-4 GHz frequency range. The amplitude measured by the SDR was not calibrated. The acquisition band, which spanned 3685 MHz to 3715 MHz, covers the transition between the 3550-3700 MHz Citizens Broadband Radio Service (CBRS) band [34] and the 3700-3980 MHz band [35], which are both used for mobile wireless communications in the United States.
@@
A spectrogram computed from the I/Q capture is shown in Fig. 6 (Top), with the frequency range cropped to 3690 MHz - 3710 MHz for this example. The spectrogram was calculated with the ‘signal.ShortTimeFFT.spectrogram’ method in the Python SciPy library [28], using a Hann window of length 512 and 50% window overlap, yielding a total of 512 frequency bins and 3517 time bins. To suppress the SDR local oscillator (LO) signal at 3700  MHz, the ‘constant’ detrending option was applied, which subtracted the mean from each window segment. The spectrogram, which visualizes the (uncalibrated) signal power across time and frequency shows visible mobile wireless emissions in the 3700-3710 MHz band, but no obvious emissions in the 3690-3700 band.
@@
PAPR was calculated in each spectrogram frequency bin by finding the ratio of the maximum to the mean across all $3517$ time bins and is plotted in Fig. 6 (Bottom) on a decibel scale. The mean PAPR in the $3690$ MHz - $3700$ MHz band was estimated to be $9.48$ dB (95% confidence interval $[9.38,9.58]$ dB) by averaging PAPR across the 170 frequency bins covering that band. The mean PAPR given by (11) with $n=3517$, was found to be $9.42$ dB, as indicated by the thick dashed line in Fig. 6 (Bottom). There was strong agreement between the observed mean PAPR in the $3690$ MHz - $3700$ MHz band and the theoretical mean value for WGN, supporting the hypothesis that the band contained only thermal noise during the observation interval. By contrast, the observed PAPR in the $3700$ MHz - $3710$ MHz band was significantly higher, consistent with mobile wireless emissions.
@@
This example illustrates how the theoretical formula (11) can be used as a quick, simple method to determine if a given frequency band contains only thermal noise, which is useful for signal detection and channel occupancy evaluations. We make three observations. First, the approach is robust to the choice of spectrogram window function. Namely, the effect of windowing cancels out in the PAPR since in the case of WGN, the value of each spectrogram frequency bin follows an exponential distribution with the distribution parameter scaled by the energy in the window sequence [32, Eq. 44]. Second, we have found this method to be fairly robust to the degree of spectrogram window overlap, e.g., 50%, 75%, although extreme degrees of windows overlap ($>90\%$) induce too much statistical dependence between window segments, rendering (11) inaccurate. Last, since PAPR is generally independent of power calibration factors, this method does not require calibrated I/Q values.
@@ VI Discussion & Conclusions
Our theoretical development was based on the assumption of white, uncorrelated I/Q samples. However, our evaluations of low-pass filtered noise (Sec. IV-C) and a spectrogram with overlapping windows (Sec. V) indicated that the expected mean PAPR is somewhat robust to data correlations. Specifically, for low-pass filtered noise with digital filter cutoff frequencies greater than $0.25$ cycles per sample, the mean PAPR was very close to the theoretical mean PAPR for uncorrelated white noise. Additionally, for a measured spectrogram, we found close agreement between the mean PAPR for thermal noise and the theoretical mean PAPR for WGN, even when the spectrogram window overlap was as high as 75%. Further investigations, both theoretical and experimental, on the impacts of data correlations on the PAPR distribution for sampled I/Q data is an interesting topic for future work.
Most of the prior literature on PAPR has focused on using it as a figure of merit for the analog RF transmitter chain in a communication system. Specifically, there has been a lot of work on PAPR reduction schemes for OFDM communication signals [1] and power amplifiers to accommodate high PAPR [2]. On the other hand, there are few references that discuss PAPR in the context of digital signal analyzer measurements, with the exception of [22, 3]. The present work helps to fill this gap by providing a simple, exact formula for the mean PAPR of WGN in I/Q data and showing how it can be applied to identify non-ideal I/Q acquisitions and to RF channel occupancy assessments.
@@ Acknowledgments
The authors thank M. Keith Forsyth, Dazhen Gu, Dan Kuester, Michelle Pirrone, and Todd Schumann for helpful comments on the manuscript.

### v3

@@ IV-C Correlated Samples
Finally, we examined the impacts of I/Q measurement correlations on the expected PAPR. Specifically, Monte Carlo simulations were performed to assess PAPR of correlated samples generated by low-pass filtering I/Q WGN. This example gives insight into the case of band-limited Gaussian noise when the (one-sided) bandwidth is smaller than the Nyquist frequency, equal to one half the sampling rate.
@@
Mean PAPR estimated from $10^{4}$ independent Monte Carlo trials for each sample size is plotted in Fig. 5 (top) together with the theoretical mean PAPR (11) for ideal I/Q WGN. In all cases, 95% confidence intervals were smaller than ±$\pm 0.025$ dB. We see that the mean PAPR generally decreased as the filter cutoff, $f_{c}$, decreased. Moreover, the mean PAPR was very close to the theoretical WGN value for $f_{c}\geq 0.25$, indicating that (11) is an accurate approximation when $0.25\leq f_{c}\leq 0.5$. Fig. 5 (bottom) compares PDFs estimated with the KDE method described in Section IV-A for the $n=10^{4}$ sample size to the I/Q WGN theoretical PDF (6).
Figure 5: Correlated noise simulations with low-pass filter cutoff frequency, $f_{c}$, specified in cycles per sample. (Top) Mean PAPR versus sample size. (Bottom) Estimated PDFs for $n=10^{4}$ samples.
@@ V Example Application to Spectral Analysis
Finally, we demonstrate how our main result, (11), can be applied to spectral analysis of sampled I/Q signals to determine if a given frequency band contains only WGN. Specifically, we illustrate the application to spectrograms, which are widely used to assess the frequency content of a nonstationary signal as it varies in time. A spectrogram is the magnitude squared of a sampled short-time Fourier transform (STFT), computed by applying the discrete Fourier transform (DFT) successively to shifted, windowed signal segments [11, Sec. 15.3].
@@
To illustrate how the above theoretical observation can be applied, we used a Deepwave Digital AIR-T 7201B [32] software defined radio (SDR) to record I/Q data at a sample rate of 30 Ms/s, for 30 ms at a center frequency of 3700 MHz. A general-purpose antenna, band-pass filter, and low-noise amplifier were used as an RF front-end prior to the SDR to amplify signals in the 3 GHz to 4 GHz frequency range. The amplitude measured by the SDR was not calibrated. The acquisition band, which spanned 3685 MHz to 3715 MHz, covers the transition between the 3550 MHz to 3700 MHz Citizens Broadband Radio Service (CBRS) band [33] and the 3700 MHz to 3980 MHz band [34], which are both used for mobile wireless communications in the United States.
@@
A spectrogram computed from the I/Q capture is shown in Fig. 6 (top), with the frequency range cropped to 3690 MHz to 3710 MHz for this example. The spectrogram was calculated with the ‘signal.ShortTimeFFT.spectrogram’ method in the Python SciPy library [27], using a Hann window of length 512 and 50% window overlap, yielding a total of 512 frequency bins and 3517 time bins. To suppress the SDR local oscillator (LO) signal at 3700 MHz, the ‘constant’ detrending option was applied, which subtracted the mean from each window segment. The spectrogram, which visualizes the (uncalibrated) signal power across time and frequency shows visible mobile wireless emissions in the 3700 MHz to 3710 MHz band, but no obvious emissions in the 3690 MHz to 3700 band.
@@
PAPR was calculated in each spectrogram frequency bin by finding the ratio of the maximum to the mean across all $3517$ time bins and is plotted in Fig. 6 (bottom) on a decibel scale. The mean PAPR in the $3690$ MHz to $3700$ MHz band was estimated to be $9.48$ dB (95% confidence interval $[9.38,9.58]$ dB) by averaging PAPR across the 170 frequency bins covering that band. The mean PAPR given by (11) with $n=3517$, was found to be $9.42$ dB, as indicated by the thick dashed line in Fig. 6 (bottom). There was strong agreement between the observed mean PAPR in the $3690$ MHz to $3700$ MHz band and the theoretical mean value for WGN, supporting the hypothesis that the band contained only thermal noise during the observation interval. By contrast, the observed PAPR in the $3700$ MHz to $3710$ MHz band was significantly higher, consistent with mobile wireless emissions.
@@
This example illustrates how the theoretical formula (11) can be used as a quick, simple method to determine if a given frequency band contains only thermal noise, which is useful for signal detection and channel occupancy evaluations. We make three observations. First, the approach is robust to the choice of spectrogram window function. Namely, the effect of windowing cancels out in the PAPR since in the case of WGN, the value of each spectrogram frequency bin follows an exponential distribution with the distribution parameter scaled by the energy in the window sequence [31, Eq. 44]. Second, we have found this method to be fairly robust to the degree of spectrogram window overlap, e.g., 50%, 75%, although extreme degrees of windows overlap ($>90\%$) induce too much statistical dependence between window segments, rendering (11) inaccurate. Last, since PAPR is generally independent of power calibration factors, this method does not require calibrated I/Q values.
@@ VI Discussion & Conclusions
Our theoretical development was based on the assumption of white, uncorrelated I/Q samples. However, our evaluations of low-pass filtered noise (see Section IV-C) and a spectrogram with overlapping windows (see Section V) indicated that the expected mean PAPR is somewhat robust to data correlations. Specifically, for low-pass filtered noise with digital filter cutoff frequencies greater than $0.25$ cycles per sample, the mean PAPR was very close to the theoretical mean PAPR for uncorrelated white noise. Additionally, for a measured spectrogram, we found close agreement between the mean PAPR for thermal noise and the theoretical mean PAPR for WGN, even when the spectrogram window overlap was as high as 75%. Further investigations, both theoretical and experimental, on the impacts of data correlations on the PAPR distribution for sampled I/Q data is an interesting topic for future work.
Most of the prior literature on PAPR has focused on using it as a figure of merit for the analog RF transmitter chain in a communication system. Specifically, there has been a lot of work on PAPR reduction schemes for OFDM communication signals [1] and power amplifiers to accommodate high PAPR [2]. On the other hand, there are few references that discuss PAPR in the context of digital signal analyzer measurements, with the exception of [3] and [21]. The present work helps to fill this gap by providing a simple, exact formula for the mean PAPR of WGN in I/Q data and showing how it can be applied to identify non-ideal I/Q acquisitions and to RF channel occupancy assessments.
@@ Acknowledgments
The authors would like to thank M. Keith Forsyth, Dazhen Gu, Dan Kuester, Michelle Pirrone, and Todd Schumann for their helpful comments on this article.